# Perception Without Vision for Trajectory Prediction: Ego Vehicle Dynamics as Scene Representation for Efficient Active Learning in Autonomous Driving

### v1

@@ II Related Research
Machine learning relies on transforming data into separable representations, and to do so effectively, requires training data which approximately covers the variance of data expected to be encountered in deployment in the real-world. To this end, active learning is a method by which data is incrementally annotated and added to a training pool for a machine learning system, selected in a strategic manner for efficiency over an annotation budget. Broadly, these methods are divided into uncertainty-driven methods, which take into account a model’s level of confidence in its prediction of an unlabeled datum, and diversity-driven methods, which take into account the relationship of a datum to all other data [13, 14]. Active learning has been useful in supporting a variety of autonomous driving tasks such as vehicle detection, recognition, and tracking [15, 16, 17].
@@ III Novelty-Sensitive Active Learning Algorithm using Trajectories and Dynamic States
Trajectory and dynamic information [23, 24] is particularly low-cost to collect and annotate. Assuming a well-calibrated GPS and IMU system, the vehicle is localized and trajectories can be reconstructed in a 2D overhead projection, along with state variables such as velocity, accleration, and heading. This requires virtually no annotation, as opposed to 2D or 3D objects in a scene, which require meticulous annotation. In the next sections, we describe ways that the low-cost information can be leveraged to curate only particular, learning-efficient scenes (which can then be expensively annotated) for an overall reduction in data budget while maintaining performance.
@@ V Analysis of Results
TABLE I: Performance of the best-performing active learning strategy in comparison to a random baseline on two common trajectory prediction metrics, taken at five data pool sizes.
Labeled Pool mADE5 mADE10 % Random Active $\alpha,\beta$ Random Active $\alpha,\beta$ 10% 1.59 1.58 (--0.01) (0%, 20%) 1.18 1.17 (--0.01) (0%, 60%) 20% 1.44 1.42 (--0.02) (20%, 20%) 1.08 1.06 (--0.02) (20%, 20%) 30% 1.37 1.35 (--0.02) (80%, 20%) 1.04 1.01 (--0.03) (60%, 20%) 40% 1.35 1.30 (--0.05) (40%, 60%) 1.00 0.97 (--0.03) (100%, 40%) 50% 1.31 1.29 (--0.02) (100%, 40%) 0.97 0.96 (--0.01) (20%, 40%) SoA 100% [28] 1.30 1.00
@@ VI Concluding Remarks
Acknowledgements
The authors would like to acknowledge the support of Qualcomm through the Qualcomm Innovation Fellowship, and thank LISA colleagues and mentors for their support.

### v2

@@ II Related Research
Machine learning relies on transforming data into separable representations, and to do so effectively, requires training data which approximately covers the variance of data expected to be encountered in deployment in the real-world. To this end, active learning is a method by which data is incrementally annotated and added to a training pool for a machine learning system, selected in a strategic manner for efficiency over an annotation budget. Broadly, these methods are divided into uncertainty-driven methods, which take into account a model’s level of confidence in its prediction of an unlabeled datum, and diversity-driven methods, which take into account the relationship of a datum to all other data [13, 14]. Active learning has been useful in supporting a variety of autonomous driving tasks such as vehicle detection, recognition, and tracking [15, 16].
@@ III Novelty-Sensitive Active Learning Algorithm using Trajectories and Dynamic States
Trajectory and dynamic information [22] is particularly low-cost to collect and annotate. Assuming a well-calibrated GPS and IMU system, the vehicle is localized and trajectories can be reconstructed in a 2D overhead projection, along with state variables such as velocity, accleration, and heading. This requires virtually no annotation, as opposed to 2D or 3D objects in a scene, which require meticulous annotation. In the next sections, we describe ways that the low-cost information can be leveraged to curate only particular, learning-efficient scenes (which can then be expensively annotated) for an overall reduction in data budget while maintaining performance.
@@ V Analysis of Results
TABLE I: Active learning strategy performance compared to random baseline at five data pool sizes.
Labeled Pool $mADE_{5}$ $mADE_{10}$ Random Active $\alpha,\beta$ Random Active $\alpha,\beta$ 10% 1.59 1.58 (--0.01) (0%, 20%) 1.18 1.17 (--0.01) (0%, 60%) 20% 1.44 1.42 (--0.02) (20%, 20%) 1.08 1.06 (--0.02) (20%, 20%) 30% 1.37 1.35 (--0.02) (80%, 20%) 1.04 1.01 (--0.03) (60%, 20%) 40% 1.35 1.30 (--0.05) (40%, 60%) 1.00 0.97 (--0.03) (100%, 40%) 50% 1.31 1.29 (--0.02) (100%, 40%) 0.97 0.96 (--0.01) (20%, 40%) SoA 100% [26] 1.30 1.00
Figure 6: From left to right: map input, predicted trajectories, and ground truth. From top to bottom, results from models trained on: 10% training data, 50% training data randomly selected, and 50% training data selected using our active learning algorithm. The 10% data model shows a large spread of possible trajectories, with little scene conformity. Though the scene conformity improves at 50% data, with active learning, the trajectories adapt even better to the lane contours of the scene.
Figure 7: As in the previous example, though the scene conformity improves at 50% data, with active learning, the trajectories snap more closely to the lanes in the freeway, while maintaining the multimodal options appropriate for the driver’s choice in the scene.
Figure 8: At the roundabout, many trajectories proposed by the 10% training data model are non-compliant, and while the 50% randomly-selected training data model shows a better conformity to the two possible modes, the ‘right turn’ mode still has four very distinct (incorrect) variants. At 50% active-learning-selected training data, these four variants collapse to one (scene-appropriate) turn.
Figure 9: Though the vehicle continues straight in this example, a good trajectory prediction model should maintain an awareness of the possibility of a left turn, as the future (in this case) is truly unpredictable since the driver has agency to elect to take the turn. The 10% training data model has no real understanding of the map, and only produces some kinematically-possible modes which are inappropriate to the scene. The 50% randomly selected training data model loses the mode for the left turn, while the active learning training method maintains both the mode and strong lane adherence in all predictions.
Qualitative results are depicted in Figures 6-9, with examples of model results from the ten percent data pool, fifty percent random data pool, and fifty percent active learning pool. Specific cases are discussed in the figure captions, and one pattern that emerges between examples is the pace of the model’s learning of lane-conforming behavior and multimodality. In the ten percent data volume, the model begins with a wide spread of modal coverage, but rarely conforming to any of the scene input. Rather, the model very loosely approximates a variety of kinematically-feasible spreads, regardless of map state. At fifty percent data randomly selected, the model begins to converge toward lane configurations, but notably deviates from the lane centerlines at a greater rate than the fifty percent data efficiently sampled using our active learning algorithm. Further, as shown in Figure 9, the active learning approach seems to show the same level of conformity even in its multi-trajectory predictions in the case of multiple possible futures.